OpenStudy (anonymous):

log with base 2(x^2-2x+5)=2 solve the equation. (The base 2 is the little 2 bottom right of log)

OpenStudy (dumbcow):

rewrite in exponential form \[x^{2}-2x+5 = 2^{2}\]

OpenStudy (anonymous):

nvm i think the answer is 1...

OpenStudy (dumbcow):

yep :)

OpenStudy (anonymous):

=]

OpenStudy (anonymous):

actually how do I work out the problem? I used a calculator, lol

OpenStudy (anonymous):

quadratic equation, or factoring, gives you 1.

OpenStudy (anonymous):

log(x^2-2x+5)/log(2)=2 is all i have

OpenStudy (anonymous):

what do i do from there?

OpenStudy (anonymous):

i think you did your change of base formula wrong the log(2) in the denominator should be log base 2 and the log in the numerator should be log base 10. However that is the hard way.

OpenStudy (anonymous):

look at the first post by dumb cow

OpenStudy (anonymous):

i just figured it out with quadratic equation :P I forgot to minus 4 on each side.

OpenStudy (anonymous):

well i meant factoring.

OpenStudy (anonymous):

\[\log_{2}(x^2-2x+5)=2 \] from here you raise both sides by the base \[2^{\log_{2}(x^2-2x+5)} = 2^{2} \] this erase the log and the equation can be simplified to \[x^2-2x+5=2^2\]